OpenStudy (anonymous):

Mutiply: 1) x^2+8x+15/x^2+3x x x^2/2x+10 2)x^2-9/x^2+x-12 x x^2-16

OpenStudy (anonymous):

\[1) \frac{ x^{2}+8x+15 }{ x^{2}+3x} \times \frac{ x^{2} }{ 2x+10 }\]

OpenStudy (anonymous):

\[2)\frac{ x^{2}-9 }{x^{2}+x-12 } \times x^{2}\]

OpenStudy (anonymous):

first, factor everything.

OpenStudy (anonymous):

So I factor what adds to what and multiples to what?

OpenStudy (anonymous):

factor the first part of 1)

OpenStudy (anonymous):

Yes I understand, but do I factor what adds to 8 and multiples to 15? I'm brainwashed right now @_@.

OpenStudy (anonymous):

And if that's correct, it would be 5 & 3.

OpenStudy (anonymous):

yes, you do. factor *everything*

OpenStudy (anonymous):

Okay so the first line would look like x(x+3)(x+5)?

OpenStudy (anonymous):

x^2(x+3)(x+5)

OpenStudy (anonymous):

Okay, how do I factor below the x^2 + 3x?

OpenStudy (anonymous):

x(x+3)

OpenStudy (anonymous):

and you can factor a 2 out of the denominator of the other fraction

OpenStudy (anonymous):

How would that look? er sorry lol .

OpenStudy (anonymous):

2(x+5)

OpenStudy (anonymous):

\[\frac{ x^{2}(x+3)(x+5) }{x(x+3)} \times \frac{ x^{2} }{ 2{(x+5)} }\]

OpenStudy (anonymous):

So then, the x+3 cancels out.

OpenStudy (anonymous):

Do the x squared's cancel out & the x+5 or not?

OpenStudy (anonymous):



OpenStudy (anonymous):

then you cancel

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x^3 / 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now the next one?

OpenStudy (anonymous):

or can you do that yourself?

hartnn (hartnn):

isn't 1st just x/2 ??

OpenStudy (anonymous):

Hartnn is gonna help me out. So I can move on to division.

hartnn (hartnn):

quite sure, 1st = x/2

OpenStudy (anonymous):

thank you hartnn, I was again making a stupid error.